OpenStudy (anonymous):

write the expression using a single exponent 2^2 x 2^5 4^10 2^10 4^16 2^16

OpenStudy (anonymous):

\[\Large a^b \times a^c =a^{b+c}\]

OpenStudy (anonymous):

Can you solve it now ?

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

a^b*a^c=a^{b+c} 2^2*2^5=2^(2+5)=2^8 Now you may figure out which of your choices has the same value of 2^8

OpenStudy (anonymous):

d